OpenStudy (imperfection):

A square and a rectangle have the same area. The length of the rectangle is 8 in. more than a side of the square, and the width of the rectangle is 4 in. less than a side of the square. Find the length of a side of the square.

mathslover (mathslover):

first of all do you know what is the area of a square?

OpenStudy (imperfection):

no

mathslover (mathslover):

area of a square = side * side and area of a rectangle = length * breadth

mathslover (mathslover):

let the side of the square = a inches

OpenStudy (imperfection):

i need to know what the length of the side of the square is.

mathslover (mathslover):

yes given that: length is 8 inches more than the side of square = 8+a right?

OpenStudy (imperfection):

ya...

mathslover (mathslover):

also : width of the rectangle is 4 inches less than the side of the square = a-4 right?

OpenStudy (imperfection):

I'm really bad at math. i need more help :(

mathslover (mathslover):

have patience i am helping you ..

OpenStudy (imperfection):

ok thank you :D

mathslover (mathslover):

now we have length of the rectangle = a+8 and width of the rectangle = a-4 area of the rectangle = length * width = (a+8)(a-4) but given area of square = area of rectangle so we have ;: \[\large{a^2=(a+8)(a-4)}\]

mathslover (mathslover):

can you tell me what is (a+8)*(a-4)

OpenStudy (imperfection):

\[a^{2}+4a-32 ?\]

mathslover (mathslover):

very good .. so we have : \[\large{a^2=a^2+4a-32}\] \[\large{\textbf{Now dubtract a^2 both sides}}\]

OpenStudy (imperfection):

4a-32?

mathslover (mathslover):

4a-32 = ?

mathslover (mathslover):

What will be in RHS there...

OpenStudy (imperfection):

a-8

mathslover (mathslover):

no no .. \[\large{\color{violet}{a^2-a^2=a^2-a^2+4a-32}}\] \[\large{\color{red}{0=4a-32}}\]

mathslover (mathslover):

add 32 both sides now

OpenStudy (imperfection):

a=8 ?

mathslover (mathslover):

right so the side of the square = 8 inches

OpenStudy (imperfection):

omg thank you so much C:

mathslover (mathslover):

best of luck for your journey here in openstudy: \[\huge{\color{red}{\cal{Welcome}}\color{green}{\space \mathbb{TO}}\color{blue}{\space \mathsf{Openstudy}}}\]